OpenStudy (anonymous):

a plane drops a package of supplies to a party of explorers. if the plane is travelling horizontally at 40m/s and is 100m above the ground where does the package strike the ground relative to the point where it released?? help plzzzz

OpenStudy (aaronq):

Once released, the package is in free-fall and its motion can be analyzed by separating the horizontal and vertical components.  In the question, you're told that the plane is moving at \(\sf v_x=100 ~m/s\), this is the horizontal component. The object is also accelerating in the vertical direction at \(\sf a_y=-9.8 ~m/s^2\) (Note the negative sign because its in the negative y direction). We don't how long that takes, but we can use the distance and the acceleration in the y direction. \(\sf \Delta y=v_{y}\Delta t +\dfrac{1}{2}a_y(\Delta t)^2\)\) \(\sf -100~m=(0)\Delta t +\dfrac{1}{2}(-9.8~m/s^2)(\Delta t)^2\) \(\sf \Delta t=4.52~s\) We can use the time for it to hit the ground with \(\sf v_x\) to find the distance it travelled. \(\sf x_f=x_i+v_x\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\) \(\sf x_f=0+(40~m/s)(4.52~s)+\dfrac{1}{2}(0)(4.52~s)^2=180.702~\approx~181~m\) The package travelled 181 meters in the horizontal direction